OpenStudy (anonymous):

through -1,3 and parallel to y=2x+1

OpenStudy (ddcamp):

Making a new line? Use the point-slope formula: \[(y-y₁) = m(x-x₁) \] where m is the slope and (x₁,y₁) is a point that it passes through. (Hint, the slope will be the same as the line it is parallel to).

OpenStudy (anonymous):

y-3=2(x+1) is that right for setting the equation up?

OpenStudy (ddcamp):

Yes, then solve for y and simplify.

OpenStudy (anonymous):

thank you. i will become a fan

OpenStudy (anonymous):

i have another question i will try and solve it first then i will ask for help.

OpenStudy (ddcamp):

ok

OpenStudy (anonymous):

through 2,2 and perpendicular to y=-3/5x+2

OpenStudy (anonymous):

i set the equation up as y-2=-5/3(x-2) but i am very bad with fractions.

OpenStudy (ddcamp):

At least with the x, you can leave it as a fraction. When you add 2 to both sides, you can rewrite it as 6/3, that should make it easier for adding the constants Also remember with perpendicular slopes, you take the negative reciprocal (flip and change the sign) for the slope.

OpenStudy (anonymous):

yes, i got to the point of y-2=-5/3x but idk how to mulitiply -5/3 and -2